Boat traveling against the current travels out 2miles, passes a log moving with the current. The boat travels 1hr more, turns around and arrives at shore the same time as the log. What is rate of current?
Answer by mananth(16949)

Let x be the speed of the rower,and y be the speed of the current.
The upstream rowing speed is x-y, mph
and the downstream rowing speed is x+y. mph
Let A be the point where the rower crosses the log.
In 1 hour after crossing the log, the rower reached x-y miles. (upstream speed*1)
The log reached y miles from A. (current speed * 1)
They are separated by distance x miles now. (x-y+y) miles =x miles
They are travelling in the same direction so subtract the speeds (x+y-y) .=x mph is the effective speed
t=d/r
x/x = 1 hr for the rower back to reach the log at the starting point.
floating 2 mile for log=====>1+1= 2hrs.
Therefore,the speed of current is 2 miles/hr.
m.ananth@hotmail.ca